The number of outcomes in the sample space for the gender of four children in a family is?The number of outcomes in the sample space for the gender of two children in a family is?
Answer #1

Possible gender = Boy or girl so 2 ways.

Number of outcomes when four children = 2*2*2*2 = 16

Number of outcomes when two children = 2*2 = 4

Please comment if any doubt. Thank you.